OpenStudy (anonymous):

solve for y... 2/9(4y - 2)=4

OpenStudy (lynfran):

first step multiply throughout by 9 ...to both sides of the equation...what do we have,..

OpenStudy (anonymous):

but you can't divide 9 and 4

OpenStudy (lynfran):

\[\frac{ 2 }{ 9 }(4y-2)=4\]the equation stay like this right?

OpenStudy (anonymous):

its 2/9 as a fraction yes

OpenStudy (lynfran):

good multiply both sides of the equation by 9

OpenStudy (lynfran):

we have 2(4y-2)=4*9

OpenStudy (lynfran):

note the 9 cancels frm the left side of the equation

OpenStudy (anonymous):

(4y -2) = 36?

OpenStudy (lynfran):

no 2(4y-2)=36

OpenStudy (anonymous):

oh ok

OpenStudy (lynfran):

now multiply parentheses

OpenStudy (lynfran):

2*4y-2*2=36

OpenStudy (anonymous):

8y - 4 = 36?

OpenStudy (lynfran):

yes

OpenStudy (anonymous):

oh ok i think i have it now

OpenStudy (lynfran):

now add 4 to both side

OpenStudy (lynfran):

oh ok

OpenStudy (anonymous):

y = 5?

OpenStudy (lynfran):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (lynfran):

welcome